### v1

# Post-Incorporating Code Structural Knowledge into Pretrained Models via ICL for Code Translation

Abstract
Code translation migrates codebases across programming languages. Recently, large language models (LLMs) have achieved significant advancements in software mining. However, handling the syntactic structure of source code remains a challenge. Classic syntax-aware methods depend on intricate model architectures and loss functions, rendering their integration into LLM training resource-intensive. This paper employs in-context learning (ICL), which directly integrates task exemplars into the input context, to post-incorporate code structural knowledge into pre-trained LLMs. We revisit exemplar selection in ICL from an information-theoretic perspective, proposing that list-wise selection based on information coverage is more precise and general objective than traditional methods based on combine similarity and diversity. To address the challenges of quantifying information coverage, we introduce a surrogate measure, Coverage of Abstract Syntax Tree (CAST), measuring maximum subtree coverage between ASTs of test source code and exemplars. Furthermore, we formulate the NP-hard CAST maximization for exemplar selection and prove that it is a standard submodular maximization problem. Therefore, we propose a greedy algorithm for CAST submodular maximization, which theoretically guarantees a $(1-1/e)$-approximate solution in polynomial time complexity. Our method is the first training-free and model-agnostic approach to post-incorporate code structural knowledge into existing LLMs at test time. Experimental results show that our method significantly improves LLMs performance in code translation and reveals two meaningful insights: 1) Code structural knowledge can be effectively post-incorporated into pre-trained LLMs during inference, despite being overlooked during training; 2) Scaling up model size or training data does not lead to the emergence of code structural knowledge, underscoring the necessity of explicitly considering code syntactic structure.
Index Terms:
Code Translation, Syntactic Validity, Large Language Model, In-context Learning.
publicationid: pubid: $0$–$00$ 20xx IEEE
I Introduction
Code translation migrates codebases from one programming language to another, avoiding the time- and labor-intensive process of redeveloping software from scratch. Automated code translation significantly enhances coding productivity in various practical scenarios, such as: 1) modernizing legacy systems (e.g., COBOL or FORTRAN) with contemporary languages (e.g., Python or Java) to extend functionality; 2) ensuring cross-platform compatibility or using language-specific advantages, e.g., integrating high-performance C++ code with Python-based machine learning frameworks.
Recently, large language models (LLMs) have demonstrated exceptional performance across various software mining tasks [1, 2, 3, 4]. However, effectively handling the syntactic structure of source code remains a significant challenge. Unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more complex, as it directly reflects program behavior. Previous studies have shown that explicitly incorporating code syntactic knowledge can significantly enhance performance in software mining tasks [5, 6, 7, 8, 9, 10, 11, 12]. However, traditional methods often depend on complex model architectures and customized loss functions, making their integration with LLMs both time-consuming and resource-intensive, even infeasible for proprietary models.
In-context learning (ICL) provides a effective and efficient solution of integrating new domain-specific knowledge into off-the-shelf LLMs without additional training by directly incorporating task exemplars into the input context [13]. For instance, in Python-to-C++ code translation, ICL enhances LLMs performance by incorporating a few paired exemplars (e.g., Python-to-C++ translation pairs) into the input prompt. This training-free approach offers flexibility across diverse tasks, avoids catastrophic forgetting, and ensures efficient resource utilization. Selecting appropriate contextual exemplars is critical for ICL. However, research on exemplar selection strategies for ICL in code translation is still limited.
Traditional ICL exemplar selection strategies often overlook code translation tasks and fail to consider the critical syntactic information inherent in source code [13]. These methods primarily rely on sample-wise similarity [14, 15, 16], ignoring compositional relationships among exemplars, which often leads to redundancy. Furthermore, syntax-aware similarity remains underexplored. Some approaches attempt to ensure diversity through clustering and additional regularization losses [17, 18, 19, 20, 21]. However, these methods often hinder the optimization of relevance and overlook compositional relationships in the syntactic structure of code. Moreover, balancing similarity and diversity poses a significant challenge.
This paper revisits exemplar selection for ICL from an information-theoretic perspective, emphasizing that information coverage serves as a more precise and general objective than the traditional similarity and diversity. However, directly quantifying information coverage in a combinatorial manner using information theory is challenging. To address this, we propose a surrogate measure for information coverage in code translation, based on Abstract Syntax Tree (AST), called the Coverage of AST (CAST). For a given test source code and exemplar set, we extract ASTs of them while retaining only node type attributes and discarding value attributes. The exemplar ASTs are then used to cover subtrees of the test source code AST. CAST quantifies the proportion of nodes in the test source code AST included in the maximum covered subtrees. Maximizing CAST supports list-wise compositional exemplar selection, achieving “two birds with one stone”: it explicitly considers syntactic structure knowledge in ICL and approximates the information coverage between the context and the test source code. However, two new challenges arise: 1) How to efficiently perform subtree covering across trees, and 2) How to solve CAST maximization for exemplar selection, as it is an NP-hard problem that cannot be optimally solved within polynomial time.
To address these challenges, we register a unique fingerprint to each AST subtree. Fingerprints for all subtrees in an AST are efficiently extracted via a single post-order traversal, and the exemplar database can be pre-processed efficiently. Next, we construct a co-occurrence matrix $M$ between the test source code and candidate exemplars, where $M_{ij}$ denotes whether the $j$-th AST subtree of the source code exists in the $i$-th candidate exemplar. Consequently, exemplar selection can be formulated as selecting rows from the co-occurrence matrix to maximize the $\ell_{1}$-norm of the OR vector across selected rows. We prove this is a standard submodular maximization problem, which has been widely studied in subset selection area. Although this is a general NP-hard problem, we can employ a greedy algorithm that starts with an empty selection set and iteratively adds elements to maximize the marginal gain of CAST at each step until the size limit $k$ is reached. This algorithm theoretically guarantees a $(1-1/e)$-approximate solution with polynomial time complexity, i.e., its performance is at least $(1-1/e)$ of the optimal solution.
To evaluate CAST, we integrate it as a plug-and-play module with state-of-the-art (SOTA) LLMs of varying scales during inference and compare it with traditional exemplar selection strategies based on similarity and diversity. Experimental results demonstrate that our method significantly improves performance over baseline methods and outperforms other exemplar selection strategies.
Our main contributions are summarized as follows:
  1. 1.
    Revealing meaningful insights: This paper presents two key findings: First, code syntactic knowledge, often overlooked during training, can be effectively post-incorporated into pre-trained LLMs at test time. Second, scaling model size or training data alone does not ensure the sufficient emergence of code syntactic knowledge, emphasizing the need to explicitly consider code syntactic structures into LLMs.
  2. 2.
    New objective for ICL exemplar selection, CAST: We propose the first training-free and LLM-agnostic approach to incorporate code syntactic knowledge into existing LLMs during inference using ICL. We revisit exemplar selection in ICL from an information-theoretic perspective, showing that information coverage is a more precise and general objective. We introduce a surrogate measure that incorporates syntactic structure knowledge into LLMs compositionally while approximating the quantify-challenging information coverage efficiently.
  3. 3.
    Practical solution for CAST maximization: Although CAST maximization in exemplar selection is NP-hard, we prove it is a submodular maximization problem and propose a greedy algorithm that theoretically guarantees a $(1-1/e)$-approximate solution with practical polynomial time complexity. Experimental results show that CAST significantly improves the code translation performance of LLMs across various scales.
II Revisiting Exemplar Selection via
Information Theory for ICL in Code Translation
II-A Preliminaries of Code Translation and ICL
Code translation can be formulated as $\mathcal{P}_{\text{trans}}{}:X\mapsto Y$, where $X$ represents the source code to be translated, $Y$ represents the target code, and $\mathcal{P}_{\text{trans}}{}$ denotes a translation model, such as an LLM. For a given source code instance $X_{\text{test}}$, the translation is formally expressed as:
$\widehat{Y}_{\text{test}}=\mathcal{P}_{\text{trans}}(\widehat{Y}_{\text{test}}|X_{\text{test}})\,.$ (1)
In ICL, the database $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ contains $n$ source and target code pairs. The indices of the database are denoted by $\mathbb{N}_{n}=\{1,2,\dots,n\}$, and $S\subseteq\mathbb{N}_{n}$ represents a subset of these indices. Thus, ICL-based code translation with $k$ exemplars as context is expressed as:
$\widehat{Y}_{\text{test}}=\mathcal{P}_{\text{trans}}(\widehat{Y}_{\text{test}}|\underbrace{X_{S[1]},Y_{S[1]},\dots,X_{S[k]},Y_{S[k]}}_{\text{context of $k$ paired exemplars}},X_{\text{test}})\,,$ (2)
where $S[i]$ represents the index of the $i$-th selected exemplar.
Therefore, the primary challenge is to select an effective context for the input source code $X_{\text{test}}$. This requires finding a subset of indices $S$ that meets the context size constraint $|S|=k$, defined as:
$S^{*}=\arg\max_{S\subseteq\mathbb{N}_{n},|S|=k}\mathrm{Score}(S,X_{\text{test}})\,.$ (3)
Traditional ICL exemplar selection relies on two main criteria: similarity and diversity [13]. Similarity-based methods select the top-$k$ most similar exemplars [22, 23, 24, 25, 14], which are simple and fast but often yield homogeneous exemplars and sub-optimal results. To address redundancy, some studies incorporate diversity [17, 18, 19], but these methods typically depend on instance-wise distance metrics and neglect the compositional relationships among exemplars. Moreover, current ICL exemplar selection methods often struggle to balance similarity and diversity, as the trade-off is task-sensitive, limiting their robustness and generalizability.
Figure 1: Revisiting ICL exemplar selection from an information-theoretic perspective; CAST for code translation.
II-B Revisiting Exemplar Selection with Information Theory
We revisit exemplar selection in ICL from an information-theoretic perspective. In code translation, information coverage refers to the overlapping information between the source code of test sample and task exemplars, quantifying the useful knowledge provided by exemplars for translating the current source code. Formally, information coverage is defined as:
$\mathrm{IC}(S,X_{\text{test}})=H(X_{\text{test}})-H(X_{\text{test}}|X_{S[1]},\dots,X_{S[k]})\,,$ (4)
where $H()$ represents information entropy. For instance, as shown in Fig. 1 (a), if two exemplars ($k=2$) with source codes $X^{{}^{\prime}}$ and $X^{{}^{\prime\prime}}$ are selected for the test source code $X_{\text{test}}$, the information coverage is:
$\mathrm{IC}(S,X_{\text{test}})=H(X_{\text{test}})-H(X_{\text{test}}|X^{{}^{\prime}},X^{{}^{\prime\prime}})$ (5)
$\quad=I(X_{\text{test}};X^{{}^{\prime}}|X^{{}^{\prime\prime}})+I(X_{\text{test}};X^{{}^{\prime\prime}}|X^{{}^{\prime}})+I(X_{\text{test}};X^{{}^{\prime}};X^{{}^{\prime\prime}})\,,$
where $I()$ denotes general mutual information, such as interaction information or conditional mutual information. Here, the required knowledge to translate $X_{\text{test}}$ ($H(X_{\text{test}})$) is divided into four parts: $I(X_{\text{test}};X^{{}^{\prime}}|X^{{}^{\prime\prime}})$ and $I(X_{\text{test}};X^{{}^{\prime\prime}}|X^{{}^{\prime}})$ measure the individual contributions of $X^{{}^{\prime}}$ and $X^{{}^{\prime\prime}}$, respectively. $I(X_{\text{test}};X^{{}^{\prime}};X^{{}^{\prime\prime}})$ represents the common knowledge provided by both $X^{{}^{\prime}}$ and $X^{{}^{\prime\prime}}$, and $H(X_{\text{test}}|X^{{}^{\prime}};X^{{}^{\prime\prime}})$ reflects the knowledge not provided by these exemplars.
Revisiting the traditional notions of similarity and diversity with the information theory, similarity can be interpreted as:
$\mathrm{Similarity}(S,X_{\text{test}})\propto\sum_{i=1}^{n}I(X_{\text{test}};X_{S[i]})\,,$ (6)
which measures the total mutual information between the test sample $X_{\text{test}}$ and each selected exemplar $X_{S[i]}$. On the other hand, pairwise diversity can be interpreted as:
$\mathrm{Diversity}(S,X_{\text{test}})\propto-\sum_{i,j=1;i\neq j}^{n}I(X_{S[i]};X_{S[j]})\,,$ (7)
which accounts for the negative mutual information between selected exemplars, ensuring reduced redundancy among them.
The Venn diagram examples in Fig. 1 illustrate the failure cases of maximizing similarity or diversity individually. Focusing solely on similarity overemphasizes common knowledge among exemplars (as shown in Fig. 1 (b)), resulting in redundant selections with minimal information gain. Conversely, optimizing diversity alone ignores the relevance between the test sample and selected exemplars and unjustly penalizes irrelevant mutual information $I(X^{{}^{\prime}};X^{{}^{\prime\prime}}|X_{\text{test}})$ (yellow part in Fig. 1 (c)). Some methods arbitrarily combine similarity and diversity as a weighted objective. To compare it with information coverage (Eq. 5), consider selecting two exemplars, $X^{{}^{\prime}}$ and $X^{{}^{\prime\prime}}$. The combined objective, weighted by $\lambda$, is given by:
$\mathrm{Similarity}(S,X_{\text{test}})+\lambda\cdot\mathrm{Diversity}(S,X_{\text{test}})$
$\qquad\propto I(X_{\text{test}};X^{{}^{\prime}})+I(X_{\text{test}};X^{{}^{\prime\prime}})-\lambda I(X^{{}^{\prime}};X^{{}^{\prime\prime}})$
$\qquad=I(X_{\text{test}};X^{{}^{\prime}}|X^{{}^{\prime\prime}})+I(X_{\text{test}};X^{{}^{\prime\prime}}|X^{{}^{\prime}})$ (8)
$\qquad\quad+(2-\lambda)I(X_{\text{test}};X^{{}^{\prime}};X^{{}^{\prime\prime}})-\lambda I(X^{{}^{\prime}};X^{{}^{\prime\prime}}|X_{\text{test}})\,.$
It is similar to Eq. 5, but there are two key differences: 1) It arbitrarily assigns different weights to the components of required knowledge. While setting $\lambda=1$ ensures fairness, increasing the selection size $k$ complicates partitioning and weighting, making the balance harder to maintain. 2) The term $I(X^{{}^{\prime}};X^{{}^{\prime\prime}}|X_{\text{test}})$ is unnecessary, as it represents shared knowledge of $X^{{}^{\prime}}$ and $X^{{}^{\prime\prime}}$. Minimizing this term is redundant and can interfere with optimizing other valuable terms.
Consequently, information coverage serves as a more precise and general objective than similarity and diversity. However, directly computing it, as defined in Eq. 4, is challenging due to the complexity of estimating the conditional distribution $\mathcal{P}(X_{\text{test}}|X_{S[1]},\dots,X_{S[k]})$. To overcome this challenge, we propose a surrogate measure called CAST, which is computationally efficient and considers code syntactic knowledge. The details are elaborated in the next section (Sec. III).
III Proposed Method
To enhance code translation in LLMs via ICL, we propose CAST, a surrogate measure for information coverage, and present a practical solution for NP-hard CAST maximization with exemplar set selection. In Sec. III-A, we define CAST, a list-wise $\mathrm{Score}$ function based on the code’s AST. CAST accounts for both the relevance of exemplars to the source code and the compositional relationships among multiple selected exemplars. In Sec. III-B, we present a greedy algorithm for CAST maximization, proving that it guarantees a $(1-1/e)$-approximation for the NP-hard subset selection problem with polynomial time complexity. Finally, Sec. III-C provides an overview of the CAST framework.
III-A Surrogate Scoring Function: CAST
We aim to integrate code structural knowledge into LLMs via ICL. As discussed in Sec. II, effective ICL exemplar selection relies on a well-designed $\mathrm{Score}$ function, with information coverage as a precise and general scoring objective. However, directly quantifying information coverage is challenging, and limited research has explored considering code structural knowledge in ICL. We propose CAST, a surrogate measure that approximates information coverage by quantifying the ratio of the maximum subtrees covered by exemplars in the test source code AST. Maximizing CAST is the first training-free and model-agnostic method for explicitly post-incorporating code structural knowledge into LLMs at test time.
We employ tree-sitter [^1] https://tree-sitter.github.io/tree-sitter to extract the AST of the source code. Each node contains two attributes: type and value, e.g.,
[…]
The type attribute reveals syntactic information, while the value attribute captures lexical details. To improve generalization, we retain only the type attribute (i.e., syntactic information) and discard the value attribute. Let $T_{\text{test}}$ denote the AST of the current test source code ($X_{\text{test}}$), and let $T_{i}$ denote the AST of the $i$-th sample’s source code ($X_{i}$) in the database. CAST serves as the $\mathrm{Score}$ function, a surrogate measure for information coverage with definition: A maximum covered subtree is a subtree of $T_{\text{test}}$ that exists in the AST of exemplars and contains no smaller covered subtrees. CAST is computed as the ratio of nodes in the maximum covered subtrees to the total nodes in $T_{\text{test}}$. The number of nodes in a tree equals the total number of its subtrees. Consequently, let $\mathrm{Sub}(T_{S})=\mathrm{Sub}(\bigcup_{i=1}^{|S|}T_{S[i]})$ denote the subtrees of ASTs for all exemplars in the context. CAST is then defined as:
$\mathrm{CAST}(S,X_{\text{test}})=\frac{\left|\mathrm{Sub}(T_{\text{test}})\cap\mathrm{Sub}(T_{S})\right|}{\left|\mathrm{Sub}(T_{\text{test}})\right|}\,.$ (10)
Input: A tree $T$ and its root node $r$.
Output: The fingerprint of the tree $T$.
1 $\mathrm{fp}\leftarrow 0$;
2 for $c\in\mathrm{children}(r)$ do
3       $\mathrm{fp}\leftarrow\mathrm{hash}(\mathrm{fp}+\mathrm{Fingerprint}(T,c))$;
4 end for
[…]
6 return $\mathrm{fp}$;
Algorithm 1 $\mathrm{Fingerprint}(T,r)$
To compute $\mathrm{CAST}$ efficiently, we assign a fingerprint ($\mathrm{fp}$) for each subtree via hashing, as detailed in Algorithm 1. The fingerprint encodes the node type (line 5) and recursively calculates the subtree fingerprint using post-order traversal and hashing. The set of fingerprints for the subtrees of a tree $T_{i}$ is denoted as $\Phi_{i}=\{\mathrm{fp}_{j}^{(i)}\}_{j=1}^{|T_{i}|}$, where all $\mathrm{fp}_{j}^{(i)}$ are computed from line 5 in Algorithm 1 by calling $\mathrm{Fingerprint}(T,r)$ only once. The fingerprint set $\Phi_{i}$ for all samples $X_{i}\in\mathcal{D}$ in the database can be precomputed. To translate a new source code $X_{\text{test}}$, we first compute the fingerprint set for subtrees in its AST $\Phi_{\text{test}}=\{\mathrm{fp}_{j}^{(\text{test})}\}_{j=1}^{|T_{\text{test}}|}$, and then construct a co-occurrence matrix $M\in\{0,1\}^{n\times|\Phi_{\text{test}}|}$, with each element $M_{ij}$ defined as:
$M_{ij}=\begin{cases}1,&\text{if }\mathrm{fp}_{j}^{(\text{test})}\in\Phi_{i}\\ 0,&\text{otherwise}\end{cases}\,.$ (11)
This matrix indicates whether the $j$-th subtree of the test source code’s AST ($T_{\text{test}}$) is present in the AST of the $i$-th source code sample ($T_{i}$) in the database.
To apply CAST as the $\mathrm{Score}$ function for list-wise exemplar selection in ICL (Eq. 2), we need to maximize CAST (Eq. 10), which is equivalent to maximizing $|\mathrm{Sub}(T_{\text{test}})\cap\mathrm{Sub}(T_{S})|$. Based on the definition of $M$, CAST maximization is formally expressed as:
$\max_{S\subseteq\mathbb{N}_{n}}$ $\quad\left\|\bigvee_{s\in S}M[s,:]\right\|_{1}$ (12)
s.t. $\quad|S|=k\,.$
Here, $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm of a vector, which counts the number of ones in the union vector $\bigvee_{s\in S}M[s,:]\in\{0,1\}^{|\Phi_{\text{test}}|}$.
This optimization problem is a subset selection task with a list-wise compositional objective, which is NP-hard that cannot be optimally resolve in polynomial time complexity.
III-B Practical Solution: Greedy Submodular Maximization
To address the NP-hard challenge of CAST maximization, we propose a greedy submodular maximization algorithm. In exemplar selection for ICL, the law of diminishing marginal utility states that as more additional exemplars are selected, the marginal (incremental) information gain is non-increasing. This property holds for both information coverage and its surrogate measure, CAST. Formally, this property is known as submodularity in mathematics and is defined as:
Definition 1 (Submodularity [26]).
Let $\Omega$ denote a finite set, and let $f:2^{\Omega}\rightarrow\mathbb{R}_{\geq 0}$ be a set function, where $2^{\Omega}$ represents the power set of $\Omega$. The function $f$ is defined as submodular if it satisfies: For every $A\subseteq B\subseteq\Omega$, and for all $x\in\Omega\setminus B$,
$f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B)\,,$ (13)
Additionally, $f$ is monotone if and only if $f(A)\leq f(B)$ for all $A\subseteq B\subseteq\Omega$.
Maximizing a non-negative, monotone, and submodular function in subset selection referred to submodular maximization. Although generally NP-hard, a greedy algorithm that iteratively maximizes the marginal gain offers a theoretical $(1-1/e)$-approximate solution in polynomial time. The surrogate CAST (Eq. 10) corresponds to the value function $f(S)=\left|\bigvee_{s\in S}M[s,:]\right|_{1}$ in the optimization problem (Eq. 12). This value function is non-negative, monotone, and submodular. We present a lemma to prove these properties:
Lemma 2 (Properties of $\ell_{1}$-norm in boolean vector operations).
Let $A,B\in\{0,1\}^{m}$. Then, the following holds:
$\|B\wedge\neg A\|_{1}=\|A\vee B\|_{1}-\|A\|_{1}=\|B\|_{1}-\|B\wedge A\|_{1}\,,$ (14)
where $\|\cdot\|_{1}$ denotes the $\ell_{1}$-norm.
Proof.
The $\ell_{1}$-norm of a binary vector $A\in\{0,1\}^{m}$ counts the number of elements equal to “$1$”:
$\|A\|_{1}=\sum_{i=1}^{m}A_{i}\,.$ (15)
Boolean operations on binary vectors are performed element-wise, as follows:
$(B\wedge\neg A)_{i}$ $=B_{i}(1-A_{i})=B_{i}-(B\wedge A)_{i}\,,$ (16)
$(A\vee B)_{i}$ $=A_{i}+B_{i}-(A_{i}\wedge B_{i})=A_{i}+(B\wedge\neg A)_{i}\,.$
Therefore,
$\|B\wedge\neg A\|_{1}$ $=\sum_{i=1}^{m}B_{i}-(B\wedge A)_{i}=\|B\|_{1}-\|B\wedge A\|_{1}\,,$
$\|B\wedge\neg A\|_{1}$ $=\sum_{i=1}^{m}(A\vee B)_{i}-A_{i}=\|A\vee B\|_{1}-\|A\|_{1}$ (17)
Consequently, the proof is complete. ∎
Proposition 3 (The value function for $\mathrm{CAST}$ maximization is non-negative, monotone, and submodular.).
Given a binary matrix $M\in\{0,1\}^{n\times m}$, the value function for an index set $S\subseteq\mathbb{N}_{n}$,
$f(S)=\left\|\bigvee_{s\in S}M[s,:]\right\|_{1}\,,$ (18)
is non-negative, monotone, and submodular.
Proof.
The function $f(S)=\left\|\bigvee_{s\in S}M[s,:]\right\|_{1}$ computes the number of ones resulting from the bitwise OR operation applied across the rows indexed by $S$.
1) Non-negativity: Since $f(S)$ counts the number of ones in the bitwise OR operation, it is always non-negative. Thus, we have $f(S)\geq 0$ for all $S\subseteq\mathbb{N}_{n}$.
2) Monotonicity: For any subsets $A\subseteq B\subseteq\mathbb{N}_{n}$, adding more rows to the bitwise OR operation cannot decrease the count of ones. Hence, $f(A)\leq f(B)$, i.e., $f(S)$ is monotone.
3) Submodularity: Given a subset $S\subseteq\mathbb{N}_{n}$, when adding a new index $x\in\mathbb{N}_{n}\setminus S$, the marginal gain is given by:
$\Delta_{S}(x)$ $=f(S\cup\{x\})-f(S)$ (19)
$=\left\|\bigvee_{s\in S\cup\{x\}}M[s,:]\right\|_{1}-\left\|\bigvee_{x\in S}M[s,:]\right\|_{1}$
$=\left\|M[x,:]\wedge\neg\bigvee_{s\in S}M[s,:]\right\|_{1}\,.$
This follows from $\|A\vee B\|_{1}-\|A\|_{1}=\|B\wedge\neg A\|_{1}$ in Lemma 2. Hence, for any subsets $A\subseteq B\subseteq\mathbb{N}_{n}$:
$\Delta_{B}(x)$ $=\left\|M[x,:]\wedge\neg\bigvee_{s\in B}M[s,:]\right\|_{1}$
$=\left\|M[x,:]\wedge\neg\bigvee_{s\in A}M[s,:]\wedge\neg\bigvee_{s\in B\setminus A}M[s,:]\right\|_{1}$
$=\underbrace{\left\|M[x,:]\wedge\neg\bigvee_{s\in A}M[s,:]\right\|_{1}}_{\Delta_{A}(x)}$ (20)
$\quad-\underbrace{\left\|M[x,:]\wedge\neg\bigvee_{s\in A}M[s,:]\wedge\bigvee_{s\in B\setminus A}M[s,:]\right\|_{1}}_{\geq 0}\,,$
where the last equation follows from Lemma 2, which states that $\|B\wedge\neg A\|_{1}=\|B\|_{1}-\|B\wedge A\|_{1}$. So, $\Delta_{A}(x)-\Delta_{B}(x)\geq 0$, thereby the value function $f$ is submodular, $\forall A\subseteq B\subseteq\mathbb{N}_{n}$ and $\forall x\in\mathbb{N}_{n}\setminus B$, the following holds:
$f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B).$ (21)
Hence, the function $f(S)$ is non-negative, monotone, and submodular. The proof is complete. ∎
Consequently, the list-wise exemplar set selection problem in Eq. 12 is a submodular maximization problem.
Submodular maximization is generally NP-hard. However, when the value function is non-negative, monotone, and submodular, a greedy algorithm can be used. It starts with an empty set and iteratively adds the element with the maximum marginal gain ($\Delta_{S}(x)$):
$x^{*}=\arg\max_{x\in\Omega\setminus S}f(S\cup\{x\})-f(S)\,,$ (22)
continuing until $k$ elements are selected. It provides a solution within polynomial time and guarantees a near-optimal approximation, as detailed in the following theorem [26, 27]:
Theorem 4 ( $(1-1/e)$-approximation of greedy submodular maximization [26]).
For any non-negative, monotone submodular function $f$, the greedy heuristic produces a solution $\hat{S}$ of size $|\hat{S}|=k$ with the following approximation guarantee:
$f(\hat{S})$ $\geq\left(1-\left(1-\frac{1}{k}\right)^{k}\right)f(S^{*})$ (23)
$\geq\lim_{k\to+\infty}\left(1-\left(1-\frac{1}{k}\right)^{k}\right)f(S^{*})$
$=\left(1-\frac{1}{e}\right)f(S^{*})\,,$
where $S^{*}$ represents the optimal solution to the submodular maximization, defined as $S^{*}=\arg\max_{S\subseteq\Omega,|S|\leq k}f(S)$.
Therefore, we propose a greedy algorithm to solve the CAST maximization problem, guided by Eq. 22. The algorithm starts with an empty selected set, $S\leftarrow\emptyset$. In each iteration, the algorithm selects a new exemplar that maximizes the marginal gain $\Delta_{S}(f)$ of Eq. 12:
$j$ $=\arg\max_{i\in\mathbb{N}_{n}\setminus S}\left\|\bigvee_{s\in S\cup\{i\}}M[s,:]\right\|_{1}-\left\|\bigvee_{s\in S}M[s,:]\right\|_{1}\,,$ (24)
[…]
$j$ $=\arg\max_{i\in\mathbb{N}_{n}\setminus S}\left\|M[i,:]\wedge\neg\bigvee_{s\in S}M[s,:]\right\|_{1}\,.$
The selected index $j$ is added to $S$ in each iteration until the subset reaches the predefined size limit $k$. The pseudocode of this greedy CAST maximization is provided in Algorithm 2.
Input: Co-occurrence matrix $M\in\{0,1\}^{n\times|\Phi_{\text{test}}|}$ and selection capacity limit $k\ll n$.
Output: Subset of exemplar indices $S\subseteq\mathbb{N}_{n}$ such that $|S|=k$.
1 Initialize the selected set: $S\leftarrow\emptyset$;
2 Initialize the coverage mask: $c\leftarrow\mathbf{0}^{|\Phi_{\text{test}}|}$;
3 for $i\leftarrow 1$ to $k$ do
[…]
5       for $j\leftarrow 1$ to $n$ do
6             if $j\notin S$ then
                   $g\leftarrow\|M[j,:]\wedge\neg c\|_{1}$; // marginal gain
[…]
9                         $s\leftarrow j$;
10                   end if
11
12             end if
13
14       end for
15      Update coverage mask: $c\leftarrow c\vee M[s,:]$;
16       Update selected set: $S\leftarrow S\cup\{s\}$;
17
18 end for
return $S$;
Algorithm 2 Greedy CAST Submodular Maximization
As shown in Theorem 4, our CAST maximization method in Eq. 24 and Algorithm 2 provides a $(1-1/e)$-approximation, ensuring sufficient effectiveness. For instance, if $k=3$ exemplars are selected, the approximation ratio reaches at least $1-(1-1/3)^{3}\approx 70.37\%$.
The time complexity of Algorithm 2 includes three parts: $\mathcal{O}(k)$ for the loop in line 3, $\mathcal{O}(n)$ for the loop in line 5, and $\mathcal{O}(|\Phi_{\text{test}}|)$ for bitwise boolean computations in line 7. Thus, the total time complexity of the CAST algorithm is $\mathcal{O}(nk\cdot|\Phi_{\text{test}}|)$. Let $L_{X_{\text{test}}}$ denote the length of the test source code. The size of the subtree fingerprint set is upper-bounded by $|\Phi_{\text{test}}|\leq L_{X_{\text{test}}}\log(L_{X_{\text{test}}})$. Moreover, the selection size $k$ and the subtree scale $|\Phi_{\text{test}}|$ are much smaller than the database size $n$, i.e., $k\ll n$ and $|\Phi_{\text{test}}|\ll n$. As a result, the greedy CAST maximization algorithm in Algorithm 2 is computationally efficient. Furthermore, the iterations in lines 5-13 are independent, allowing the loop in line 5 to be parallelized for faster computation. With sufficient resources in industrial applications, this optimization reduces the time complexity to $\mathcal{O}(k\cdot|\Phi_{\text{test}}|)$, further improving efficiency.
Constructing the co-occurrence matrix takes $\mathcal{O}(n\cdot|\Phi_{\text{test}}|)$ time, based on its size ($M^{n\times|\Phi_{\text{test}}|)}$). Extracting the subtree fingerprint set $\Phi_{\text{test}}$ from the test source code takes $\mathcal{O}(|\Phi_{\text{test}}|)$ time using a single post-order traversal on $T_{\text{test}}$. Moreover, the subtree fingerprint sets for database samples can be precomputed, avoiding extra computations at test time. An inverted index links each subtree fingerprint $\mathrm{fp}$ to the corresponding database samples. For a new test source code, the inverted index (e.g., using ElasticSearch) efficiently retrieves co-occurrence coordinates for each $\mathrm{fp}\in\Phi_{\text{test}}$ in $\mathcal{O}(|\Phi_{\text{test}}|)$ time. Each $\mathrm{fp}$ retrieves a sparse column (in $\{0,1\}^{n}$) for the co-occurrence matrix $M\in\{0,1\}^{n\times|\Phi_{\text{test}}|}$. Combining sparse columns to form the full matrix $M$ takes $\mathcal{O}(n\cdot|\Phi_{\text{test}}|)$.
In summary, constructing the co-occurrence matrix for CAST takes $\mathcal{O}(n\cdot|\Phi_{\text{test}}|)$, while the greedy CAST algorithm in Algorithm 2 has a polynomial time complexity of $\mathcal{O}(nk\cdot|\Phi_{\text{test}}|)$. Using distributed and parallel computation, this can be further reduced to $\mathcal{O}(k\cdot|\Phi_{\text{test}}|)$. Thus, our approach ensures both theoretical effectiveness with a near-optimal $(1-1/e)$-approximation and practical efficiency.
Figure 2: Overview of post-incorporating code syntactic knowledge into pre-trained models using ICL based on CAST retrieval.
III-C Overall Framework of CAST
Figure 2 illustrates the overall pipeline of our approach, which post-incorporates code syntactic structure knowledge into off-the-shelf LLMs through ICL based on Covering Abstract Syntax Tree (CAST) retrieval. To accelerate the overall process, lexical similarity is used to coarsely pre-recall $\lfloor t\cdot k\rfloor$ candidate samples, where $t\geq 1$ is a hyperparameter. Specifically, the Levenshtein distance is employed to recall the top-$\lfloor t\cdot k\rfloor$ nearest candidates. The fingerprint set of the test source code $\Phi_{\text{test}}$ is extracted using Algorithm 1, and a co-occurrence matrix $M$ is constructed with all pre-recalled candidate samples. Then, Algorithm 2 selects $k$ exemplars as context. The selected $k$ exemplar translation pairs are directly concatenated with the test source code to create the prompt, as shown in Fig. 2. This prompt is then input into pre-trained models, such as LLMs, to translate the test source code.
Besides, our CAST retrieval method includes two variants: 1) CAST-F: Selects a fixed number of exemplars based on the predefined capacity limit $k$. 2) CAST-A: Automatically determines the number of exemplars using a CAST (defined in Eq. 10) threshold, allowing early stop when the coverage rate reaches the threshold.
IV Experiments
To evaluate CAST, we integrate it as a plug-and-play module with SOTA LLMs of varying scales during inference and compare it with traditional exemplar selection strategies based on similarity and diversity. Extensive experiments are conducted to evaluate the effectiveness and efficiency of CAST on thousands of translation tasks.
IV-A Dataset and Database
We use the dataset from [28], containing 948 parallel functions in Python, Java, and C++, widely used in code translation [28, 29, 30]. Of these, 568 include unit tests for at least one programming language: 464 for Python, 482 for Java, and 467 for C++. Following the previous work [31], our experiments focus on this subset. ICL requires a separate database. For each pre-trained LLM, we use it to translate the entire test dataset and verify the predictions using test input cases. Samples that pass all test input cases are added to the database for this LLM. Models are evaluated on the remaining data, and final metrics are computed using the entire dataset, treating the database samples as exactly correct.
IV-B Baselines
We apply CAST-F and CAST-A to SOTA LLMs and compare them with LLMs using various ICL exemplar selection strategies, as well as several learning-based translation models.
TABLE I: Results of pre-trained models, including LLMs and code translation models, with fixed exemplar strategy.
Models C++ to Python Python to C++ Java to C++ C++ to Java Average
CA EM CA EM CA EM CA EM CA EM
Off-the-shelf Large Language Models
CodeGen-6B 30.17% 4.40% 22.70% 1.94% 35.12% 7.57% 11.20% 2.46% 24.80% 4.09%
CodeLLaMA-13B 23.24% 2.29% 16.90% 2.11% 11.62% 3.35% 16.73% 3.52% 14.48% 2.82%
CodeLLaMA-34B 45.07% 6.16% 39.08% 4.23% 23.59% 7.57% 38.38% 14.96% 36.53% 8.23%
CodeLLaMA-70B 44.01% 4.58% 50.00% 3.70% 49.65% 16.37% 43.66% 11.80% 46.83% 9.11%
DeepSeek-Coder-6.7B 44.97% 7.22% 19.75% 3.14% 34.04% 13.73% 53.17% 17.78% 37.98% 10.47%
DeepSeek-Coder-33B 53.70% 7.75% 28.87% 4.01% 39.44% 14.96% 54.06% 18.31% 44.02% 11.26%
Gemma2-2B 47.36% 6.51% 39.61% 2.64% 54.05% 12.15% 22.18% 5.28% 40.80% 6.65%
Gemma2-9B 60.39% 8.45% 59.26% 4.75% 65.61% 31.34% 20.81% 23.06% 51.52% 16.90%
Gemma2-27B 63.38% 8.45% 63.91% 5.33% 68.49% 24.96% 68.31% 21.83% 66.02% 15.14%
Qwen2-7B 58.27% 6.16% 50.70% 4.40% 60.74% 13.20% 50.88% 11.62% 55.15% 8.85%
Qwen2.5-72B 69.01% 8.63% 68.49% 6.34% 71.30% 23.42% 64.61% 24.30% 68.36% 15.67%
LLaMA3-70B 69.37% 5.63% 60.74% 10.03% 67.96% 23.39% 64.61% 27.99% 65.67% 16.76%
GPT-4o-mini 70.95% 9.33% 69.01% 6.16% 70.95% 6.69% 54.58% 22.01% 66.37% 11.05%
GPT-4o 71.65% 9.49% 68.66% 5.99% 72.71% 23.02% 65.85% 26.41% 69.72% 16.23%
Learning-based Code Translation Models
TransCoder 36.64% 2.29% 30.40% 0.88% 27.84% 5.81% 49.77% 14.39% 36.16% 5.84%
TransCoder-IR / / / / 40.99% 14.26% 50.53% 18.28% 45.76% 16.27%
TransCoder-ST 46.34% 2.64% 47.85% 1.06% 49.68% 9.68% 64.73% 17.43% 52.15% 7.70%
  • $\Phi$
    For clarification, we adopt “-” to concatenate LLM and corresponding parameter size for discrimination. For example, Qwen2 with 7B parameters is dubbed Qwen2-7B.
IV-B1 Baseline Code Translation Models
We compare our method with recent LLM families across various parameter scales, consisting of 14 advanced LLMs:
  • CodeGen [2]: An LLM trained in natural language and programming data for conversation-based program synthesis. We employ CodeGen-6B-Multi in the experiment.
  • Qwen [32, 33]: A suite of models with parameter sizes ranging from 0.5B to 72B, including dense and Mixture-of-Experts variants, excelling in language, multilingual tasks, coding, math, and reasoning.
  • Gemma [34]: A lightweight LLMs, ranging in scale from 2B to 27B parameters.
  • DeepSeek-Coder [1, 35]: A series of open-source code LLMs, pre-trained on 2 trillion tokens using the fill-in-the-blank task to enhance code generation and infilling.
  • LLaMA [36, 37]: Multilingual LLMs with parameter sizes from 8B to 70B, are designed to enhance inference, code generation, and instruction compliance.
  • CodeLLaMA [3]: A series of code-centric LLMs derived from LLaMA2 [37], further trained on 500B code tokens.
  • GPT-4 [38]: Advanced generative AI models developed by OpenAI. Although not explicitly trained for code generation, they demonstrate notable performance in this domain. Their effectiveness in handling code generation tasks is primarily attributed to their massive scale.
Moreover, three state-of-the-art learning-based code translation models are served as baselines.
  • TransCoder [28]: An unsupervised code translation model using monolingual samples, pre-trained with cross-lingual modeling, denoising, and back-translation.
  • TransCoder-IR [29]: A model extended on TransCoder that incorporates compiler intermediate representation (IR) to enhance code translation.
  • TransCoder-ST [30]: A model extended on TransCoder that utilizes automatically generated test cases to filter out invalid translation results.
IV-B2 Baseline ICL Exemplar Selection Strategies
We evaluate several baseline ICL exemplar selection strategies to compare with our proposed CAST variants. These strategies encompass simple heuristic methods, similarity-based approaches, and diversity-oriented techniques, as detailed below:
  • Fixed Exemplars (Fixed): A static set of exemplars is pre-selected and consistently applied to all test samples.
  • Random Sampling (Random): Exemplars are randomly selected from the database for each test sample, serving as a baseline for comparison without any guided selection.
  • Similarity with Levenshtein Distance (LD): Exemplars are selected via lexical similarity to the test source code, measured by the Levenshtein distance.
  • Similarity on CodeBERT Embedding (CodeBERT): Exemplars are retrieved using cosine similarity scores between CodeBERT [39] embeddings of the test source code and candidate samples.
  • Diversity with Clustering (Diversity): Candidate samples are grouped into $k$ clusters by $k$-means based on the CodeBERT [39] embeddings, and the most similar sample from each cluster is selected to ensure diversity.
IV-C Evaluation Metrics
We employ Computational Accuracy (CA) and Exact Match Accuracy (EM) to evaluate model performance, following the prior works [28, 29, 30, 31].
Traditional studies primarily use the Bi-Lingual Evaluation Understudy (BLEU) [40] metric for evaluating code translation as a natural language translation task. However, in code translation, correctness of code execution is more critical. CA measures the percentage of translated code that produces identical execution results to the ground truth code under the same inputs. This metric assesses execution consistency, assuming the provided inputs adequately cover all boundary conditions, and is defined as:
$\mathrm{CA}\triangleq\mathop{\mathbb{E}}\limits_{X_{\text{test}}}\left[\mathbb{I}\left(\mathrm{Exec}(X_{\text{test}},\mathcal{D}_{X}^{\text{test}})=\mathrm{Exec}(\widehat{Y}_{\text{test}},\mathcal{D}_{X}^{\text{test}})\right)\right]$ (25)
$\mathcal{D}_{X}^{\text{test}}$ denotes the input set for the current source code $X_{\text{test}}$. $\mathrm{Exec}(X_{\text{test}},\mathcal{D}_{X}^{\text{test}})$ and $\mathrm{Exec}(\widehat{Y}_{\text{test}},\mathcal{D}_{X}^{\text{test}})$ indicate the outputs produced by executing the source code $X_{\text{test}}$ and the translated target code $\widehat{Y}_{\text{test}}$ with the input cases $\mathcal{D}_{X}^{\text{test}}$, respectively.
Additionally, since the datasets provide ground truth code in the target programming languages, we evaluate EM accuracy, defined as the proportion of target translations that exactly match the ground truth target code.
IV-D Empirical Results and Analysis
TABLE II: Results of LLMs of various parameter scales for 5-shot Code Translation upon different ICL strategies.
Models Strategy C++ to Py Py to C++ Java to C++ C++ to Java Py to Java Java to Py Average
Qwen2-7B Random 63.38% 59.31% 66.54% 61.27% 49.74% 67.43% 61.28%
LD 64.96% 57.69% 67.96% 63.73% 52.03% 67.61% 62.33%
CodeBERT 64.79% 57.02% 66.02% 63.91% 53.35% 66.37% 61.91%
Diversity 64.96% 59.51% 67.42% 65.15% 51.41% 67.61% 62.68%
CAST-F 65.32% 61.62% 68.84% 65.15% 55.77% 69.01% 64.28%
Gemma2-27B Random 67.25% 67.08% 70.77% 71.83% 66.02% 69.89% 68.81%
LD 67.43% 66.73% 70.95% 69.54% 65.49% 69.54% 68.28%
CodeBERT 67.25% 66.55% 70.42% 70.25% 64.79% 69.72% 68.16%
Diversity 68.13% 66.55% 70.77% 69.37% 64.08% 69.89% 68.13%
CAST-F 70.60% 68.66% 71.13% 72.49% 66.90% 70.40% 70.03%
Qwen2.5-72B Random 69.54% 70.07% 73.94% 66.73% 61.80% 72.01% 69.02%
LD 70.25% 70.42% 71.48% 70.95% 59.15% 72.18% 69.07%
CodeBERT 70.42% 69.72% 71.65% 70.42% 60.56% 72.01% 69.13%
Diversity 70.25% 69.89% 73.06% 71.30% 64.96% 71.65% 70.19%
CAST-F 70.77% 70.77% 75.53% 73.06% 65.31% 72.54% 71.33%
[…]
question-circle RQ1: How do SOTA LLMs of various scales perform in the code translation task?
In this research question, we evaluate SOTA LLMs across various scales for code translation. As shown in Tab. I, the results reveal several meaningful observations: 1) The LLMs demonstrate impressive performance in the code translation task, highlighting their potential for practical application. Specifically, the proprietary LLMs, such as GPT-40 and GPT-4o-mini, achieve an average CA of 66.37% and 69.72%, respectively, indicating strong potential for LLM-based code translation. 2) A high EM score indicates that the translated code is identical to the ground truth, often used as a metric to assess the extent of training data leakage. Compared to learning-based methods, which ensure no data leakage during training, the EM scores of LLMs are comparable, indicating minimal data leakage. This shows that LLMs can effectively understand the task through prompts and produce semantically correct translated code. 3) Most LLMs outperform learning-based models. Within the same model series, larger models, such as the Gemma2 series, perform better than smaller ones, indicating that scaling up parameters enhances the code translation capabilities of LLMs.
[…]
Figure 3: The performance of Qwen2-7B with various strategies and shot count.
[…]
question-circle RQ2: How does CAST perform with SOTA LLMs compared to different ICL exemplar selection strategies?
Referring to the results in Tab. I, we choose three open-source LLMs as primary baselines: Qwen2-7B, the SOTA LLM under 10B parameters; Gemma2-27B, the SOTA LLM under 40B parameters; and Qwen2.5-72B, the SOTA LLM exceeding 70B parameters. CAST-F is used to select exemplars for ICL, and its performance is compared against other exemplar selection strategies. Results in Tab. II demonstrate the effectiveness of CAST-F across LLMs with various parameter scales in the 5-shot code translation task. There are several key observations: 1) CAST-F consistently outperforms all other ICL exemplar selection strategies across LLMs and translation tasks. Specifically, it surpasses the second-best strategy, Diversity, by 1.6% for Qwen2-7B, 1.9% for Gemma2-27B, and 1.14% for Qwen2.5-72B. This shows that information coverage is a more precise and general measure for ICL, and CAST-F effectively integrates code structural knowledge into LLMs to enhance code translation. 2) For Gemma2-27B, exemplar selection using LD, CodeBERT, and Diversity performs worse than random sampling. This highlights the importance of effective exemplar selection strategies in ICL. CAST-F, the only list-wise method that considers both information coverage and code structural knowledge, consistently achieves superior results compared to other strategies.
[…]
TABLE III: Performance of Qwen2-7B on code translation with various ICL exemplar selection strategies and numbers of exemplars.
Strategy #Shots C++ to Py Py to C++ Java to C++ C++ to Java Py to Java Java to Py Average
Fixed 1 58.27% 50.70% 60.74% 50.88% 20.07% 63.20% 50.64%
Random 1 60.39% 54.93% 66.90% 52.11% 29.93% 67.60% 55.31%
3 63.02% 57.02% 65.14% 61.90% 49.74% 67.78% 60.77%
5 63.38% 59.31% 66.55% 61.27% 49.74% 67.43% 61.28%
10 63.73% 57.57% 66.55% 63.73% 49.65% 67.78% 61.50%
15 63.20% 59.86% 66.73% 64.96% 50.70% 67.42% 62.15%
20 63.56% 59.15% 66.02% 64.08% 51.94% 68.49% 62.21%
LD 1 59.51% 59.51% 67.43% 54.57% 34.51% 67.78% 57.22%
3 61.44% 60.92% 67.78% 63.38% 51.06% 67.43% 62.00%
5 64.96% 57.69% 67.96% 63.73% 52.03% 67.61% 62.33%
10 64.26% 60.73% 67.43% 65.67% 52.46% 68.49% 63.17%
15 63.73% 59.15% 66.90% 65.49% 52.99% 67.25% 62.59%
20 64.44% 59.15% 67.43% 64.43% 52.64% 67.25% 62.56%
CodeBERT 1 64.61% 56.69% 64.61% 54.92% 46.13% 67.78% 59.12%
3 63.91% 57.39% 66.73% 62.15% 47.05% 67.25% 60.75%
5 64.79% 57.02% 66.02% 63.91% 53.35% 66.37% 61.91%
10 64.79% 56.16% 65.49% 64.08% 53.17% 68.66% 62.06%
15 65.32% 55.96% 65.32% 64.43% 53.17% 67.78% 62.00%
20 64.44% 57.75% 65.49% 61.62% 54.23% 68.49% 62.00%
Diversity 1 64.61% 56.69% 64.61% 54.92% 46.13% 67.78% 59.12%
3 64.08% 58.63% 67.42% 64.26% 54.05% 67.96% 62.73%
5 64.96% 59.51% 67.42% 65.15% 51.41% 67.61% 62.68%
10 64.79% 59.33% 67.42% 65.49% 53.12% 68.31% 63.08%
15 64.96% 60.39% 67.42% 65.85% 53.52% 67.96% 63.35%
20 65.14% 60.74% 66.02% 65.49% 52.29% 68.84% 63.09%
CAST-F 1 64.08% 61.09% 67.96% 59.96% 35.72% 68.13% 59.52%
3 64.61% 61.09% 68.31% 64.44% 54.74% 68.13% 63.58%
5 65.32% 61.62% 68.84% 65.15% 55.77% 69.01% 64.28%
10 65.14% 60.92% 68.49% 66.19% 56.16% 69.01% 64.32%
15 65.84% 61.09% 68.66% 66.37% 57.22% 69.19% 64.73%
20 66.02% 61.27% 69.37% 66.55% 58.10% 69.37% 65.11%
Average #Shots 11.43 15.65 8.08 13.98 9.57 7.37 13.94 11.43
CAST-A ($\bm{\geq 98\%}$) 67.25% 61.80% 71.13% 68.31% 56.16% 69.37% 65.67%
[…]
question-circle RQ3: How does ICL perform using different exemplar selection strategies as the number of exemplars increases?
Results in Tab. III and Fig. 3 show the performance of Qwen2-7B in code translation using various ICL exemplar selection strategies, with the number of exemplars increasing from $1$-shot to $20$-shots. Several key observations arise from this analysis: 1) CAST-F consistently achieves the highest average performance across all strategies, with an accuracy of 65.11% at $20$-shots, surpassing the second-best strategy, Diversity, by 2.02%. 2) Increasing the number of exemplars generally improves performance across most strategies. CAST-F demonstrates superior scalability, particularly at higher shot counts, underscoring its robustness in improving code translation. 3) Compared to baselines, CAST-F consistently delivers incremental gains by leveraging information coverage and code structural knowledge. 4) The Diversity strategy ranks second, performing well by considering both point-wise similarity and diversity, but it lacks the ability to account for code structural knowledge or perform selection in a compositional manner. 5) CAST-F achieves the highest accuracy across challenging language pairs, such as Python to Java (58.10%) and Java to Python (69.37%), demonstrating its effectiveness in translating between syntactically diverse languages. 6) A law of diminishing marginal utility is evident in Tab. III and Fig. 3 as the number of exemplars increases. While performance improves significantly with the first few exemplars, additional gains become minimal beyond $5$-shots due to redundancy or information saturation. However, Qwen2-7B can remit this issue effectively by considering information coverage, as a more precise and general measure.
[…]
(a) C++ to Java
(b) Python to C++
(c) Python to Java
(d) C++ to Python
Figure 4: Comparison of the average coverage of abstract syntax tree (CAST) between the CAST-F and Levenshtein distance (LD) strategies.
[…]
question-circle RQ4: How does CAST-A perform when dynamically determining the selection size?
[…]
question-circle RQ5: How does the CAST convergence of CAST-F compare with that of LD?
CAST serves as a surrogate measure of information, indicating the ratio of necessary knowledge gained from exemplars during code translation. As shown in Fig. 4, CAST is computed from exemplars retrieved by CAST-F and LD across varying shot counts, revealing several key observations: 1) CAST-F consistently converges faster than LD as the number of shots increases, demonstrating greater efficiency in acquiring necessary knowledge from exemplars for translation tasks. 2) CAST-F achieves a higher final CAST ratio than LD across all translation tasks, showcasing its robustness in optimizing exemplar selection. This emphasizes the advantage of CAST-F in handling complex translations.
[…]
TABLE IV: Results of the gain of CAST-F among different sizes of LLMs, consisting of 2B, 9B, and 27B of Gemma2.
Models Strategy C++ to Python Python to C++ Java to C++ C++ to Java Average
CA EM CA EM CA EM CA EM CA EM
Gemma2-2B Fixed 47.36% 6.51% 39.61% 2.64% 54.05% 12.15% 22.18% 5.28% 40.80% 6.65%
CAST-F 51.06% 8.10% 48.77% 3.17% 63.56% 17.43% 30.81% 10.39% 48.55% 9.77%
Improv. 7.81% 24.42% 23.13% 20.08% 17.59% 43.46% 38.91% 96.78% 21.86% 46.18%
Gemma2-9B Fixed 60.39% 8.45% 59.26% 4.75% 65.61% 31.34% 20.81% 23.06% 51.52% 16.90%
CAST-F 60.74% 14.08% 64.08% 9.51% 69.19% 36.27% 65.14% 32.22% 64.79% 23.02%
Improv. 0.58% 66.63% 8.13% 100.21% 5.46% 15.73% 213.02% 39.72% 56.80% 55.57%
Gemma2-27B Fixed 63.38% 8.45% 63.91% 5.33% 68.49% 24.96% 68.31% 21.83% 66.02% 15.14%
CAST-F 70.60% 11.62% 68.66% 6.87% 71.13% 34.86% 72.49% 28.87% 70.72% 20.56%
Improv. 11.39% 37.51% 7.43% 28.89% 3.85% 39.66% 6.12% 32.25% 7.20% 34.58%
[…]
question-circle RQ6: Does scaling up the model size automatically lead to sufficient code structural knowledge?
As shown in Tab. II, although translation performance improves with larger models, the improvement from explicitly incorporating the code structural knowledge by CAST-F is much more substantial. Additionally, for a fairer comparison, we evaluate CAST-F with $5$-shots on Gemma2-(2B/9B/27B), all of which are from the same family. The results, shown in Tab. IV, indicate that the improvement from CAST-F is not significantly related to model scale. This confirms that the additional code structural knowledge from CAST-F is independent of model scale. Therefore, simply scaling up the model or training data does not result in the sufficient emergence of code structural knowledge, highlighting the necessity of explicitly incorporating code syntax.
[…]
Figure 5: Sensitivity analysis of hyper-parameters in the average CA of Qwen2-7B for the selection size $k$ and the pre-recalled candidate size $\lfloor t\cdot k\rfloor$.
Figure 6: The distribution of result types in Python-to-Java 20-shots translation by Qwen2-7B across various exemplar selection strategies in ICL. Outcomes are categorized as Compile Error, Runtime Error, Time Limit Exceeded, Wrong Answer, and Accept.
[…]
question-circle RQ7: What is the hyperparameter sensitivity?
In CAST-F, there is only one hyperparameter, i.e., the factor $t$ for adjusting the size of lexical pre-recall ($\lfloor t\times k\rfloor$). We evaluate CAST-F across varying values of the factor $t$ and the selection size $k$. Fig. 6 shows the results, which indicate that regardless of the selection size $k$, $t=2$ yields the best performance, and performance remains stable across all values of $k$ for varying $t$. For CAST-A, there is an additional threshold $\tau$. Results in Fig. 3 indicate that CAST-A achieves Pareto optimality when $\tau$ varies between 90% and 98%, demonstrating stable and effective performance across different thresholds.
[…]
question-circle RQ8: What is the impact of our method on failure frequencies?
This study evaluates how exemplar selection strategies influence test code failures across various input cases. Results in Fig. 6 highlight key observations: 1) CAST-F achieves the lowest proportion of “Wrong Answer” outcomes (26.41%) and the highest “Accept” rate (58.10%), surpassing Levenshtein Distance (52.64%) and CodeBERT Similarity (54.23%). This underscores its superior effectiveness in selecting exemplars to enhance translation accuracy. 2) ICL yields limited improvements for “Compile Error” and “Runtime Error.” These limitations likely stem from cross-language differences and the lack of comprehensive third-party library knowledge, which hinder LLMs’ ability to resolve such errors.
[…]
V Threat Analysis
Our results are interpreted with three threats to validity.
  • The internal validity threat comes from potential LLM data leakage due to possible training/testing set overlaps, but prior work [31] shows this threat is limited.
  • The first external validity threat is in the implementation of the compared techniques. To reduce it, we select state-of-the-art pre-trained models and reuse their implementations and weights. We believe this threat is minimal.
  • Another external validity threat is the quality of the dataset. It was released by [28] and is widely used due to multi-lingual samples and self-contained test suites [29, 30, 31]. We use the cleaned dataset from [31], where four authors cross-checked their cleaned results in pairs to reach a consensus for each group of parallel functions.
  • The construct threat lies in the evaluation metrics. To reduce this threat, we adopt CA and EM to evaluate the code translation performance of different approaches from both lexical and semantic correctness as most of the relevant papers [28, 29, 30, 31].
VI Related Works
VI-A Code translation
Automated code translation aids in the migration of codebases across programming languages, significantly enhancing productivity in software development. Early rule-based translation methods have been extensively explored [41, 42, 43, 44]. Furthermore, phrase-based statistical machine translation techniques have shown notable advances [45, 46, 47, 48, 49, 50]. Due to the success of deep learning, deep code translation methods have extended the SOTA further [51, 52, 53, 54, 55].
Prior research has shown that, unlike the grammatical structure of natural language, the syntactic structure of source code is stricter and more intricate, directly reflecting the program’s behavior [6, 56, 57, 58, 9]. Previous studies have demonstrated that explicitly incorporating structural knowledge into models enhances their perceptual understanding of code, thereby aiding software mining tasks [11, 12], including code translation [10, 6]. For example, modeling structural knowledge from AST [5, 6, 7], control flow graphs (CFG) [8, 9], and data flow graphs (DFG) [10]. However, these syntax-aware methods depend on meticulously crafted model architectures and loss functions tailored through training.
Recently, LLMs, scaled to hundreds of billions of parameters, have exhibited remarkable performance across various natural language processing tasks [59, 60, 61]. By training on vast code corpora, several LLMs have achieved impressive performance, such as GPT series [62, 38], LLaMA series [37, 36], Qwen series [32, 33], and Gemma series [63, 34]. Furthermore, to specifically advance in code tasks, some LLMs are pre-trained and continuously trained with more code data and code-specific tasks, including CodeLLaMA [3], CodeGemma [64], Qwen2.5-Coder [33], Starcoder series [4, 65], and DeepSeek-Coder series [1, 35]. However, mainstream LLMs often overlook code structural knowledge, whose effectiveness has been validated in prior methods. Directly combining the carefully designed model architectures and loss functions of classic syntax-aware methods with LLMs is both time- and resource-intensive, even be prohibited for proprietary models. Therefore, research on how to post-incorporate code structural knowledge into off-the-shelf LLMs is limited but crucial.
VI-B In-context Learning
Recent studies have shown that LLMs exhibit strong ICL abilities as model and data sizes scale up [66, 67, 68]. The essence of ICL lies in learning by analogy, requiring a prompt context composed of a few demonstration exemplars. By concatenating these exemplars with a query, LLMs can make better predictions without updating parameters, differentiating ICL from traditional model tuning. Recently, numerous studies have attempted to leverage ICL for complex tasks, including natural language understanding [14], reasoning [69, 70, 71], text generation [23, 72, 22]. Additionally, several studies aim to provide theoretical justifications and insights into how LLMs learn from a few in-context demonstration exemplars [73, 74]. Despite significant progress, ICL still faces challenges and requires further exploration, such as its sensitivity to the selection of demonstration exemplars [75].
Various objectives for ICL exemplar selection and tailoring have been explored [76, 77, 16, 13]. Mostly ICL exemplar selection typically relies on point-wise similarity [14, 15, 16], where a matching score is computed between each candidate exemplar and the input query, followed by the independent selection of the top-$k$ exemplars to construct the context [13]. Although these methods are straightforward and fast, they fail to consider compositional relationships among the selected $k$ exemplars, often leading to redundancy caused by homogeneity. Some studies aim to ensure diversity by clustering exemplars into $k$ groups, then selecting the most similar exemplars from each group for a given query [17, 18, 19]. Besides, some works [20, 21] introduce an auxiliary loss to regularize diversity. However, additional diversity loss may hinder the optimization of relevance, and accounting for compositional relationships in diversity remains challenging. Some methods attempt to combine weighted similarity and diversity, but adaptively balancing these two objectives is difficult. Moreover, these methods do not focus on code translation and are typically syntax-unaware, neglecting the critical syntactic information inherent in source code. Additionally, [21] attempts to account for local structure coverage in ICL. However, it only considers the structure between a parent node and its child nodes, and the algorithm is rough, needing to work together with complex similarity and diversity regularization losses. In contrast, our method is grounded in information theory, providing greater precision and generality. We consider a more comprehensive subtree structure, and our algorithm can be applied directly, guaranteed by submodular theory, without requiring additional losses.
VII Conclusion and Future Work
Code translation aids code migration between languages, boosting productivity in software development. This work focuses on post-incorporating code structural knowledge into pre-trained models via ICL for code translation. We revisit exemplar selection in ICL from an information-theoretic perspective, showing that maximizing information coverage provides a more precise and general solution than traditional methods. We introduce CAST, a surrogate measure based on the maximum subtree coverage ratio of AST. We show that NP-hard CAST maximization is a submodular maximization problem and propose a greedy algorithm with a $(1-1/e)$-approximate solution of polynomial time complexity. Our method is the first training-free, model-agnostic approach to integrate code structural knowledge into existing LLMs at test time, and both theory and experiments prove their effectiveness and efficiency. This work provides two key insights: 1) Code structural knowledge can be post-incorporated into pre-trained LLMs during inference despite training neglect. 2) Scaling model size or data doesn’t lead to code structural knowledge, highlighting the need to consider code syntax in LLMs.
Although our method shows promise in incorporating code structural knowledge into pre-trained models, several areas for future improvement remain. For instance, considering more advanced code structure beyond AST, e.g., CFG and DFG, could provide richer information for software mining in pretrained models. Besides, introducing a feedback loop during inference, where the model refines exemplar selection based on previous outputs, could further enhance performance.
Acknowledgments
This research is supported by the National Natural Science Foundation of China (NSFC) under grant numbers 62076121 and 61921006, the Major Program (JD) of Hubei Province (2023BAA024), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (KYCX24_0301).